Lauren shot a rocket from her backyard.The equation that models this scenario is y=-16x^2+80x+10. For how long was Laurens rocket above 74 feet?
Answer by TimothyLamb(4379)

---
y = -16x^2 + 80x + 10 = 74
-16x^2 + 80x + 10 = 74
-16x^2 + 80x - 64 = 0
---
the above quadratic equation is in standard form, with a=-16, b=80, and c=-64
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
-16 80 -64
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the two real roots (i.e. the two x-intercepts), of the quadratic are:
x = 1
x = 4
---
answer:
the rocket was above 74 feet altitude between 1 second and 4 seconds after launch
so for a total of 3 seconds
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php